OpenStudy (rainbow_dash):

What is the inverse of y=3x^3-3

jimthompson5910 (jim_thompson5910):

Swap x and y to get x = 3y^3 - 3 Now solve for y

OpenStudy (rainbow_dash):

so y = \[\sqrt[3]{x-3}/3\]

jimthompson5910 (jim_thompson5910):

close, it should be \[\Large y = \frac{\sqrt[3]{x+3}}{3}\] Notice is +3 and NOT -3

OpenStudy (rainbow_dash):

oh my bad simple mistake. thanks!

jimthompson5910 (jim_thompson5910):

np